OpenStudy (anonymous):

A bowl contains 8 blue marbles and 5 green marbles. Hannah randomly draws 4 marbles from the bowl. She does not replace the marbles after each draw. What is the probability that she draws 2 blue marbles and 2 green marbles?

OpenStudy (anonymous):

a.28/429 b.5/429 c.4/13 d.56/143

OpenStudy (campbell_st):

well it looks like you are selecting P(B, B, G, G) = 8/13 * 7/12 * 5/11 * 4/10

OpenStudy (anonymous):

im sorry i dont understand what you mean :/

OpenStudy (campbell_st):

well there are 13 marbles 8 blue and 5 green so if the 1st marble is selected is a blue, then P(B) = 8/13 since the marble isn't replaced there are now only 7 blue marbles and 12 in total... so selecting a 2nd blue marble is 7/12 so selecting a green next, there is 5 of them but only 11 left... the 2nd green marble is 4 from 10 so multiply th probabilities to get P(B, B, G, G) in that order

OpenStudy (anonymous):

so nthe answer would be d.56/143

OpenStudy (rational):

you need to consider other permutations also

OpenStudy (anonymous):

A?

OpenStudy (campbell_st):

but that's on the assumption that of the 17160 possible you only want B, B, G, G because if they way 2 blue and 2 green in any order...it changes... and for the solution I posted you might like to multiply the probabilities again.

OpenStudy (campbell_st):

well A is my best guess for a very badly worded question.

OpenStudy (rational):

guess we need to consider all the permutations http://www.wolframalpha.com/input/?i=%284+choose+2%29*8%2F13+*+7%2F12+*+5%2F11+*+4%2F10

OpenStudy (anonymous):

thank you very much

OpenStudy (rational):

which one is correct ?

OpenStudy (anonymous):

rational you were right :)

OpenStudy (anonymous):

8c2*5c2/13c4=answer

OpenStudy (anonymous):

280/715 further simplify it

OpenStudy (anonymous):

56/143